Work done by a substance in reversible nonflow manner in accordance with V=100/p3 ft3 where P in psi. Eavaluate the



work done on or by the substance as the pressure increases from 10 psia to 100 psia.




1
Expert's answer
2022-06-06T08:52:56-0400
"W=\int_{V_1}^{V_2}pdV"

"W=\int_{p_1}^{p_2}p(-50/p^2)dp=-50\ln \frac{p_2}{p_1}"

"=-50\ln \frac{100}{10}=-115\:\rm psia\cdot ft^3"